OpenStudy (anonymous):

What is the y-intercept of the equation 2x + 3y = 12? -2/3 4 6 12

OpenStudy (astrophysics):

Hey set it up in y = mx+b (b = y - intercept) format that way you will be able to see the y - intercept clearly.

OpenStudy (astrophysics):

So solve for y, do you know how to do that?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

do it

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

2x + 3y = 12 start by subtracting 2x both sides

OpenStudy (astrophysics):

What do you mean

OpenStudy (anonymous):

u - 2x from 12 =10 2x - 3y and u get 1y

ganeshie8 (ganeshie8):

remember this in algebra : you can't add/subtract apples and coconuts

ganeshie8 (ganeshie8):

2x + 3y = 12 subtracting 2x both sides : 2x + 3y = 12 -2x -2x you should get : 3y = 12 - 2x

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

but "y" still has that 3 attached to it how to get rid of that 3 ?

OpenStudy (astrophysics):

Often for these algebraic problems to move something to the other side we use the opposite operation of what already exists, so since there's a 3 attached to the y, which means it's being multiplied. What can we do to get rid of the 3?

OpenStudy (astrophysics):

Don't worry about being wrong, say what ever is on your mind it's ok

OpenStudy (anonymous):

c

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

i am not doing that

OpenStudy (astrophysics):

Yeah that's much better...haha divide both sides by 3

OpenStudy (astrophysics):

I've been up for 24+ hours so that's my queue xD

OpenStudy (anonymous):

i did not delete your things

OpenStudy (astrophysics):

I know I did

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

hey just one moment

OpenStudy (astrophysics):

Let me do the algebra quickly to...

ganeshie8 (ganeshie8):

4 is correct, but may i know this : when you said below, what were you thinking ? 

OpenStudy (astrophysics):

3y=12-2x y=(12-2x)/3 y= -2x/3+4 works out

OpenStudy (anonymous):

me or him

OpenStudy (astrophysics):

\(\color{blue}{\text{Originally Posted by}}\) @ltj1234 u have 10 1 y left \(\color{blue}{\text{End of Quote}}\) Ganeshie is asking you what you meant by your comment here

OpenStudy (anonymous):

o ok

ganeshie8 (ganeshie8):

yes you @ltj1234

OpenStudy (anonymous):

idk

OpenStudy (astrophysics):

The 10 I think you were trying to subtract 2x from 12, but you can't do that since there is a variable (x)

ganeshie8 (ganeshie8):

sure you know it ^

ganeshie8 (ganeshie8):

when you have `12 - 2x`, you leave it like that you can't say that equals `10` because they are not "like terms" nevertheless, it was really a very good attempt from your side! keep it up!

OpenStudy (anonymous):

ok

OpenStudy (astrophysics):

For sure, good job @ltj1234

OpenStudy (anonymous):

thanks